The ratio of the measures of the three angles of a triangle is 3:4:8. Find the measure of the largest angle.
Answer by mananth(16949)

ratio = 3:4:8
Total parts = 3+4+8=15
3/15 *180=36
4/15 *180=48
8/15 *180 = 96
..
96+48+36=180